Water rises in a capillary tube to a certain height such that the upward force due to surface tension is balanced by 75 × 10-4 N, forces due to the weight of the liquid. If the surface tension of water is
6 × 10-2 N/m, the inner-circumference of the capillary must be
(a) 1.25 × 10-2 m (b) 0.50 × 10-2 m
(c) 6.5 × 10-2 m (d) 12.5 × 10-2 m

Grade:12

1 Answers

Daniyal
7 years ago
The upward force due to the change in pressure in the capillary.F(up)=∆P*A =(2*S.T.*cos(angle of contact)/R)*πR^2 assuming angle of contact to be zero =2*π*R*S.T.=(circumference)*(6*10^(-2)) This F(up)=75*10^(-4) From there you will get Circumference is 12.5*10^(-2)m.